OpenStudy (anonymous):

\[1^{st}\text{ order ODE }\\ (xy+1)dx+(2y-x)dy=0\]

OpenStudy (anonymous):

Working with exact equations?

OpenStudy (anonymous):

yes but this one is not exact and the intergration factor is ugly

OpenStudy (anonymous):

\[M(x,y)=xy+1~~\Rightarrow~~\frac{\partial M}{\partial y}=x\\ N(x,y)=2y-x~~\Rightarrow~~\frac{\partial N}{\partial x}=-1\] Our integrating factor can be either \[\exp\left(\int\frac{M_y-N_x}{N}~dx\right)~~\text{or}~~\exp\left(\int\frac{N_x-M_y}{M}~dy\right)\] \[\int\frac{x+1}{2y-x}~dx~~\text{or}~~-\int\frac{x+1}{xy+1}~dy\] The second integral looks like it might be easier to work with, so I'd try that.

OpenStudy (anonymous):

but our proffesor told us we shud not do intergration factor with two variables...like this one,....

OpenStudy (anonymous):

so i was thinking maybe another way cud do cos we didnt learnt multivariate function

OpenStudy (amistre64):

is it seperable with a y=vx substitution?

OpenStudy (anonymous):

when i did that substitution it gives, \[\large (vx^2+2v^2x-vx+1)dx+x^2(2v-1)dv=0\] still not exact,dint c hw this cn be seperable

OpenStudy (anonymous):

very intrasting one i'm also trying .........